### v2

# DESI results and Dark Energy from QCD topological sectors

Abstract
[…]
show]waerbeke@phas.ubc.ca
I Introduction
Since the discovery of the accelerated expansion rate of the Universe interpreted as a mysterious dark energy (DE) Perlmutter et al. (1999); Riess et al. (1998) a countless number of models have been proposed to explain its existence. The recent DESI results are the most precise measurement of the DE equation of state so far DESI Collaboration et al. (2025). If these results are confirmed, explaining simultaneously $w_{\rm DE}>-1$ today and a redshift evolution such that $w_{\rm DE}(z>0)<-1$ will represent a real challenge for DE models.
The possibility that $w_{\rm DE}$ may vary with redshift came as a surprise, although there are theoretical arguments suggesting that such a dependence is to be expected Brandenberger (2025). Even more unexpected is the possibility that $w_{\rm DE}$ could cross the $w_{\rm DE}=-1$ line. The regime where $w_{\rm DE}<-1$—commonly referred to as the phantom regime—is already highly puzzling, as it leads to violations of unitarity and causality when formulated within the framework of quantum field theory (QFT). The additional feature of crossing the $w_{\rm DE}=-1$ boundary—known as quintom behavior—further challenges fundamental principles of QFT, as discussed in the review by Cai et al. (2010). For a recent summary of this topic, including numerous original references and context in light of the DESI results, we refer the reader to the short overview in Cai et al. (2025). The origin of all the issues associated with the $w_{\rm DE}\leq-1$ behavior lies in the presence of a dynamical canonical field $\phi$ governed by a potential $V(\phi)$. In such cases, the theory cannot be formulated consistently due to internal instabilities and related problems, including a negative square of the speed of sound, $c_{s}^{2}<0$, among other fundamental inconsistencies.
In this paper, we develop a dramatically different approach based on decade old idea, proposed by Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018), which finds its root in a 1967 paper by Zeldovich Zeldovich (1967). Essentially, the Zeldovich’s prescription (expressed in modern terms) can be formulated as follows. One must compute the non-perturbative vacuum energy in Quantum Chromodynamics (QCD) in an expanding universe, characterized by dimensional parameters such as the QCD scale $\Lambda_{\mathrm{QCD}}$ and the Hubble constant $H$. A similar computation should then be performed for the QCD vacuum energy in Minkowski spacetime. The difference between these two results—referred to as the leftover in this work—enters the right-hand side of the Einstein equation, in accordance with Zeldovich’s prescription.
A common objection to this proposal is that QCD, which describes all known nuclear physics, involves very short-range forces and therefore should not be sensitive to a parameter like the Hubble constant. However, a central point of this proposal is that this is not true for certain specific observables—most notably, the vacuum energy and the associated contact terms. In a strongly coupled gauge theory with multiple topological sectors, these terms can be sensitive to arbitrarily large distances. This insight opens the possibility for a time-dependent dark energy component, governed by the modification of these topological sectors and the tunnelling transition rates between them in a time-dependent, expanding background characterized by the Hubble parameter $H(t)$. In this approach, there are no new degrees of freedom such as a scalar field $\phi$, and therefore none of the issues associated with the phantom or quintom regimes arise. These problems simply do not emerge in our framework, as the entire construction is rooted in Standard Model (SM) physics, without introducing any new fields—albeit in a highly nontrivial way. All relevant scales in this framework are determined by the leftovers resulting from Zeldovich’s subtraction prescription.
The paper is organized as follows. In Section II, the central idea is introduced in three steps. First, Section II.1 develops an intuitive picture of QCD-induced dark energy. This is followed by Sections II.2 and II.3, where we present the core conceptual results of the framework—both in terms of amplitude and timing (i.e., why it becomes relevant now)—expressed through a single, well-known dimensional parameter of QCD, $\Lambda_{\mathrm{QCD}}$, without introducing any new fields or coupling constants. In Section III, we derive the solution to the Friedmann equation within our framework and analyze it in both the asymptotic limit $t\rightarrow\infty$ and for redshifts $z\geq 0$. We then discuss several possible cosmological implications of our approach and how it can be tested, before concluding in Section IV.
II The topology as the source of the gravitating vacuum energy
The objective of this section is to provide an overview of the basic ideas proposed in Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018) concerning the nature of vacuum energy in strongly coupled gauge theories such as quantum chromodynamics (QCD). In this framework, the Universe will naturally evolve towards the de Sitter geometry, as a result of the presence of topologically nontrivial sectors $|k\rangle$ in a time-dependent, expanding universe In our proposal, the Dark Energy (DE) is identified with the vacuum energy produced by tunnelling transitions between different topological sectors. The amplitude of DE is associated with the QCD energy scale, while its time variation arises from changes in the topological sectors $|k\rangle$ and the corresponding tunnelling transition rates in an expanding universe.
It may be instructive to develop an intuitive picture—presented in Sect. II.1—of the vacuum energy in this framework, formulated through an analogy with condensed matter (CM) systems. This will illustrate, with a simple picture, why the topology of the vacuum in gauge theories is important. Such an intuitive perspective can be helpful in gaining a basic understanding of the nature of DE within the framework proposed in this work.
It is important to emphasize that many key elements of this proposal have been tested in simplified gauge theories, where the computations can be done analytically. The relevant references, along with a list of important technical components used in this construction, are provided in subsections II.2 and II.3.
II.1 Intuitive picture
The intuitive picture arises from the fact that quantized gauge fields can give rise to non-local effects of a topological nature. We will argue in Section II.2 that DE could be the manifestation of such non-local, long range, effect taking place in the present time.
In principle, gauge theories with a mass gap 111this corresponds to theories where all excitations are massive and there are no physical massless fields. Physically, QCD exhibits confinement, and there are no massless free gluons observed in nature. The lightest physical excitations (e.g. pions) have non-zero mass, implying a mass gap between the vacuum and the lowest energy state. are expected to have short-range interactions due to exponential suppression of correlators at large distances (as per cluster decomposition). However, several non-local or topological effects in gauge theories can give rise to long-range forces or correlations, even in the presence of a mass gap. We propose that the observed DE in the Universe may be the manifestation of such a non-local effect.
Consider, for example, the Aharonov-Casher effect Reznik and Aharonov (1989): an external charge is inserted into a superconductor where the electric field is exponentially screened, $\sim\exp(-r/\lambda)$, with $\lambda$ being the penetration depth. Despite the very short penetration length, a neutral magnetic fluxon remains sensitive to the external charge even at arbitrarily large distances $r\gg\lambda$, despite the screening of the physical field (which is equivalent to the presence of a mass gap in our system). This genuine quantum effect is purely topological and non-local in nature, and it can be understood in terms of the dynamics of the gauge sectors responsible for the long-range behavior.
Now imagine studying the same effect in a time-dependent background—an analogy for the expanding Universe. The corresponding topological sectors $|k\rangle$ will be modified due to the variation of the external background. However, this modification cannot be described in terms of any local dynamical fields, since there are no propagating long-range fields in the system since the physical electric field is screened. By analogy with the Aharonov-Casher effect, our proposal is that the time evolution of Dark Energy (DE) is governed by the modification of the topological sectors $|k\rangle$ as the Universe expands.
We conclude this discussion on the nature of Dark Energy (DE)—and its sensitivity to arbitrarily large distances despite the presence of a mass gap in QCD—by noting that the existence of long-range forces in gauge theories with a gap has been suspected for a long time. This idea appears, for instance, in a paper by Lüscher Luscher (1978) titled ”The secret long range force in quantum field theories with instantons.” The technical origin of these long-range forces can be traced back to an earlier work by ’t Hooft ’t Hooft (1976), where the computation of the instanton density requires introducing an infrared (IR) cutoff to properly account for zero gauge modes. A crucial aspect of this calculation is that the correction due to the IR cutoff is only power suppressed—not exponentially suppressed, as one might naively expect from the presence of a gap in the system. See also Thomas and Zhitnitsky (2012) for several relevant comments on this point. The same long-range interaction can be described in terms of the so-called Veneziano ghost Veneziano (1979); see Appendix A. In exactly solvable models, one can explicitly see that the Veneziano ghost is in fact an auxiliary, topological, non-propagating field—commonly used in condensed matter physics to describe topologically ordered systems—which lacks a canonical kinetic term. We briefly review the connections between these various descriptions of this highly nontrivial phenomenon with specific references in Appendix A.
II.2 Generating DE from QCD topological sectors
In the approach of Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018), the vacuum energy entering the Friedmann equation is defined as $\Delta\rho\equiv\rho_{\rm FRW}-\rho_{\mathrm{Mink}}$, where $\rho_{\mathrm{Mink}}$ represents the vacuum energy in Minkowski spacetime. This definition of vacuum energy was first proposed in 1967 by Zeldovich Zeldovich (1967), who argued that $\rho_{\text{vac}}=\Delta\rho\sim Gm_{p}^{6}$ must be proportional to the gravitational constant, with $m_{p}$ being the proton mass (i.e., the QCD scale). In the following decades, various papers, using the same definition for $\Delta\rho$ in Einstein’s field equations, have been written by researchers in different fields, including particle physics, cosmology, condensed matter physics, see Zhitnitsky (2015) for the references and details.
The computation of $\Delta\rho$ for arbitrary geometries is currently infeasible due to several technical challenges (see Zhitnitsky (2015) for a detailed discussion). However, in certain special cases, such calculations have been successfully performed and form a key part of the theoretical framework developed in Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018). One such case involves the relativistic hyperbolic spacetime $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$, characterized by the curvature parameter $\kappa$, where $\mathbb{H}^{3}_{\kappa}$ denotes three-dimensional hyperbolic space, and $\mathbb{S}^{1}_{\kappa^{-1}}$ is a one-dimensional circle (or ring) with circumference $\kappa^{-1}$. The vacuum energy associated with this geometry is denoted by $E_{\rm vac}[\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}]$. In comparison, the vacuum energy of flat Minkowski spacetime is given by $E_{\rm vac}[\mathbb{R}^{3}\times\mathbb{S}^{1}]\sim\Lambda_{\rm QCD}^{4}$, where $\mathbb{R}^{3}$ is three-dimensional Euclidean space and $\mathbb{S}^{1}$ is a one-dimensional circle. According to the Zeldovich prescription as described above, the central claim of Zhitnitsky (2015) is that the vacuum energy difference between these two geometries, $\Delta E_{\rm vac}$, receives a linear correction proportional to $\kappa$, i.e.:
$\Delta E_{\rm vac}\equiv E_{\rm vac}[\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}]-E_{\rm vac}[\mathbb{R}^{3}\times\mathbb{S}^{1}]\approx$ (1)
$-\left[\Lambda_{\mathrm{QCD}}^{4}\left(1-c_{\kappa}\frac{\kappa}{\Lambda_{\mathrm{QCD}}}\right)-\Lambda_{\mathrm{QCD}}^{4}\right]\approx c_{\kappa}\kappa\cdot\Lambda_{\mathrm{QCD}}^{3},$
where numerical factors are omitted and $c_{\kappa}$ is a dimensionless numerical coefficient of order one. The crucial minus sign ()(-) is retained, which shows that the vacuum energy difference is positive. The negative sign associated with the QCD vacuum energy is a well-known feature of QCD, originally introduced through the so-called Bag constant, a phenomenological parameter in the MIT bag model Chodos et al. (1974a, b). This Bag constant was later shown to be expressible in terms of the energy-momentum tensor and its trace anomaly. The negative sign, proportional to $c_{\kappa}\kappa$, arises from explicit computations Zhitnitsky (2015), and can be interpreted as the negative sign typically encountered in Casimir-type energy calculations for systems with boundary conditions or constraints, in contrast to Minkowski space-time.
For de Sitter spacetime, it is not currently possible to perform analogous computations Zhitnitsky (2015). One of the main obstacles is that monopole solutions with nontrivial holonomy contributing to $\Delta E_{\rm vac}$ —similar to those defined on $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$— are not known in the context of de Sitter spacetime. However, as argued in Barvinsky and Zhitnitsky (2018), one can conjecture that the resulting expression is expected to closely resemble Eq.(1). It means that, in de Sitter spacetime, holonomy is expected to emerge dynamically (as the Universe expands), and the role of $\kappa$ in Eq.(1) is assumed by the Hubble parameter in the de Sitter Universe, with the replacement $c_{\kappa}\sim 1$ by $c_{H}\sim 1$.
Therefore, in this framework, DE is induced by the tunneling between QCD vacuum topological sectors in an expanding Universe, leading to an emerging positive DE density given by $\rho_{\rm DE}=\Delta E_{\rm vac}\propto H$ in vicinity of the de Sitter state. We refer to Appendix B with more technical details of our QCD induced DE proposal, here we list a few important physical properties:
1. All effects discussed in this work are non-perturbative in nature and non-analytic in the QCD coupling constant, scaling as $\propto\exp(-1/g^{2})$. As such, they cannot be captured within QCD perturbation theory  [^2] Non-perturbative computations in QCD are typically carried out in $\mathbb{R}^{4}$ space with a Euclidean signature, where tunneling transitions are described using classical solutions such as instantons, calorons, and similar configurations. These solutions are usually defined in Euclidean space, after which an analytic continuation is performed to translate the results into physical space-time with a Lorentzian signature..
2. Furthermore, all effects discussed here are global in nature and cannot be formulated in terms of any local effective QFT. This is fundamentally different from the conventional treatment where dark energy is described as a new fundamental scalar field $\phi$ (quintessence”, K-essence”, “phantom fields”, etc.) and an effective potential $V(\phi)$. These approaches suffer from numerous issues, including the fine-tuning problem, instabilities, violations of unitarity, and other critical principles of quantum field theory. By contrast, in our QCD-induced DE mechanism, there are no new dynamical degrees of freedom (as explained intuitively in Section II.1). As a result, our framework is free from any violations of the fundamental principles mentioned above.
3. The relevant topological Euclidean configurations that saturate the vacuum energy (1) can be interpreted as three-dimensional magnetic monopoles wrapping around the $\mathbb{S}^{1}$ direction Zhitnitsky (2015). These configurations are characterized by non-vanishing holonomy —a measure of the gauge field’s behavior around a closed loop—, which ultimately leads to a linear (rather than quadratic) correction $\sim\kappa$ to the vacuum energy density. Additional discussion of this linear dependence can be found in Appendix B with references on the original results. This entire gauge configuration represents only a saddle point in the Euclidean path integral—a mathematical construction used to describe quantum tunnelling events. It does not correspond to the propagation of a physical degree of freedom capable of transmitting information or signals. In a cosmological context, such configurations are highly unconventional: they inherently describe non-local physics, since the holonomy is itself a non-local quantity. This non-locality is precisely why these effects cannot be captured by any local quantum field theory, as noted in item 2. It is a similar situation to the Aharonov-Casher effect, which was discussed in Section II.1 to provide an intuitive picture.
4. Equation (1) is consistent with earlier findings in the weakly coupled “deformed QCD” model, where all computations are analytical. In that model, the sensitivity of the vacuum energy to very large distances can be studied by enclosing the system in a box of size $\mathbb{L}$. As shown in Thomas and Zhitnitsky (2012), the corrections to the vacuum energy scale linearly with the inverse size, $\sim\mathbb{L}^{-1}$—a behavior analogous to the role played by the parameter $\kappa$ in equation (1). This model closely resembles the system considered in the present work, since the vacuum energy in the deformed QCD model is also dominated by monopoles with nontrivial holonomy and features a mass gap. In contrast, conventional ’t Hooft instantons with trivial holonomy produce only quadratic corrections, $\sim\mathbb{L}^{-2}$, as also noted in Thomas and Zhitnitsky (2012).
5. Equation (1) is also consistent with lattice simulations presented in Yamamoto (2014), where the author investigates the rate of particle production in a de Sitter background. The results show that the production rate is linearly proportional to the Hubble constant, scaling as $\sim H$, rather than the expected $H^{2}$.
In our framework, the de Sitter behavior in Lorentzian spacetime is not driven by a local, dynamical dark energy field $\phi$. The driving mechanism can be interpreted as a Casimir-type vacuum energy [^3] This novel form of Casimir energy is a genuine physical effect that could, in principle, be tested through tabletop experiments, as suggested in the conclusion., arising from numerous tunnelling transitions in a strongly coupled gauge theory. This energy is determined by the QCD scale parameter $\Lambda_{\rm QCD}$ and is characterized by the ratio ${\Lambda_{\mathrm{QCD}}^{3}}/{M_{PL}^{2}}$. In this context, our framework replaces the dimensional parameters typically introduced via the dark energy potential $V(\phi)$ in standard cosmological models with this QCD-derived quantity.
II.3 The Equation of State (EoS) for QCD-induced DE in de Sitter case
Based on the conjecture that Eq. (1) can be applied in a de Sitter universe by replacing $\kappa$ with the Hubble parameter, we can estimate the order of magnitude of the dark energy density today, assuming the Universe is approaching a de Sitter phase—a scenario supported by the dominance of the observed dark energy density parameter. To this end, we introduce the notation $\overline{H}$, rather than the observed value $H$, to emphasize that Eq. (1) pertains to the asymptotic value of the Hubble constant, attained when the Universe approaches a de Sitter phase with a scale factor evolving as $a(t)\propto\exp(\overline{H}t)$. In this asymptotic regime, the dark energy and the Hubble constant $\overline{H}$ acquire the following values:
$\overline{H}^{2}=\frac{8\pi G}{3}\rho_{\rm DE},~~\rho_{\rm DE}$ $\approx$ $c_{H}\Lambda_{\mathrm{QCD}}^{3}\overline{H},~~G\equiv M_{PL}^{-2}$
$\overline{H}=c_{H}\frac{8\pi\Lambda_{\mathrm{QCD}}^{3}}{3M_{PL}^{2}},~\rho_{\rm DE}$ $\approx$ $c_{H}^{2}\frac{8\pi\Lambda_{\mathrm{QCD}}^{6}}{3M_{PL}^{2}},$ (2)
This corresponds precisely to the estimate originally proposed by Zeldovich long ago Zeldovich (1967), provided one replaces $m_{p}\rightarrow{\Lambda_{\mathrm{QCD}}}$ in his formula. Taking $\Lambda_{\mathrm{QCD}}\approx 100\;{\rm MeV}$ and assuming that the dimensionless numerical coefficient $c_{H}$ remains constant [^4] We do not lose any generality by fixing $\Lambda_{\mathrm{QCD}}\approx 100\;{\rm MeV}$. This is because the dimensionless parameter $c_{H}$ can always be redefined to absorb all numerical coefficients that arise in the calculations. In particular, a small numerical QCD-related factor $\propto m_{q}/\Lambda_{\mathrm{QCD}}\approx 0.05$, which consistently accompanies tunnelling transitions in QCD, is also absorbed into $c_{H}$. In this estimate, the quark mass is taken to be approximately $m_{q}\approx 5$ MeV.:
$\overline{H}$ $=$ $c_{H}\frac{8\pi\Lambda_{\mathrm{QCD}}^{3}}{3M_{PL}^{2}}\approx\bar{c}_{{H}}\cdot 2.8\cdot 10^{-33}eV,$ (3)
$t_{0}$ $\equiv$ ${\overline{H}}^{-1}=\frac{7.3}{\bar{c}_{{H}}}{\rm Gyr},~~c_{H}\equiv\bar{c}_{{H}}\left(\frac{m_{q}}{\Lambda_{\mathrm{QCD}}}\right)\approx 0.05\bar{c}_{{H}},$
$\rho_{\rm DE}$ $=$ $c_{H}\Lambda_{\mathrm{QCD}}^{3}\overline{H}\approx\bar{c}_{{H}}^{2}\left(3.4\cdot 10^{-3}{\rm eV}\right)^{4},$
which are indeed very close to the observed values today [^5] It is also worth mentioning that the numerical coincidence between the observed value of $\rho_{\rm DE}$ and the estimate in equation (3) was the primary motivation for the proposal in Urban and Zhitnitsky (2010a, b), which suggested that the driving force behind dark energy is the nontrivial dynamics of the topological sectors in strongly coupled QCD—originally formulated in terms of the Veneziano ghost (albeit without a clear understanding at the time of the physical basis for the formula). It took several years before the key elements outlined in items 1–5 of Section II.2 were fully understood. The connection between the Veneziano ghost and the topological auxiliary field was also elucidated later in Zhitnitsky (2013). Additional remarks on this relationship are provided in Appendix A..
In the conventional treatment of dark energy, where models involve a field $\phi$ and its potential $V(\phi)$, the values of $\rho_{\rm DE}$ and $t_{0}$ are typically introduced in an ad hoc fashion. In contrast, within our framework, the order of magnitude of these parameters emerges naturally and is determined solely by a single QCD scale, $\Lambda_{\mathrm{QCD}}$. This approach has the potential to address several fine-tuning problems, including the “coincidence problem” (why is dark energy relevant now?), the “drastic separation of scales,” and the “unnatural weakness of interactions.”
It then becomes relatively straightforward to compute the corresponding DE Equation of State (EoS) using standard thermodynamic principles:
$dF=TdS-PdV,~~~~P=-\left.\frac{\partial F}{\partial V}\right|_{S},$ (4)
to arrive
$P$ $=$ $-\frac{\partial F}{\partial V}=+\Lambda_{\mathrm{QCD}}^{4}\left(1-c_{H}\frac{\overline{H}}{\Lambda_{\mathrm{QCD}}}\right)$
$\rho$ $=$ $\frac{F}{V}=-\Lambda_{\mathrm{QCD}}^{4}\left(1-c_{H}\frac{\overline{H}}{\Lambda_{\mathrm{QCD}}}\right).$ (5)
The resulting subtraction procedure—removing a large contribution from the energy and pressure associated with flat Minkowski spacetime, as discussed above, leads to the following expressions for the vacuum pressure and energy in an expanding Universe:
$P_{DE}=-c_{H}\Lambda_{\mathrm{QCD}}^{3}\overline{H},~~~\rho_{DE}=+c_{H}\Lambda_{\mathrm{QCD}}^{3}\overline{H}$ (6)
such that the equation of state assumes the form
$w$ $=$ $\frac{P_{DE}}{\rho_{DE}}=-1,~~{a}(t)\propto\exp(\overline{H}t),$ (7)
which is precisely the behaviour describing the de Sitter Universe.
The main results of this subsection can be summarized as follows: The pure de Sitter state can be characterized by a single parameter—the Hubble constant $\overline{H}$—and the tunnelling transitions in QCD generate dark energy with the same characteristics as a cosmological constant, as given by (6) and its equation of state (7). We emphasize that no ad hoc constants are introduced into the system. In principle, the constant $c_{H}$ could be computed from first principles (although this is not currently feasible, as discussed in Zhitnitsky (2015)), similar to model computations briefly reviewed in Sect. II.2. This constant is dynamically generated, as Zeldovich conjectured more than half a century ago Zeldovich (1967), if one replaces $m_{p}\rightarrow\Lambda_{\mathrm{QCD}}$ in his formula. What happens if we slowly vary H{H} over time? According to the adiabatic theorem, the relations (6) and (7) are expected to hold as long as the de Sitter state continues to dominate the cosmic evolution. We will make use of this feature in the next Sect. III, where we investigate deviations from exact de Sitter behavior (7) with the aim of comparing to recent DESI results.
III DESI results and deviation from exact de Sitter behaviour
The DESI results DESI Collaboration et al. (2025) suggest not only that the dark energy equation-of-state parameter $w$ differs from 1-1, but also that it may vary over time and even cross the w=1w=-1 boundary. If confirmed, such behavior would present significant challenges for conventional field theory models. The aim of this section is to examine whether this kind of behavior can arise within our proposed framework. To do so, we begin by representing the Friedmann equation in the context of our approach and analyzing its solutions.
III.1 The Friedmann’s equation with QCD-induced dark energy
The Friedmann equation is:
$1=\Omega_{\rm m}+\Omega_{\rm r}+\Omega_{\rm DE}$ (8)
where $\Omega_{\rm m}(z)$, $\Omega_{\rm r}(z)$, $\Omega_{\rm DE}(z)$ are the matter, radiation and DE density parameters at redshift $z$. We assume a zero curvature Universe. We then rewrite the density parameters as a function of the scale factor $a\propto(1+z)^{-1}$ and the Hubble parameter $H(z)$:
$\Omega_{\rm m}$ $=$ $\frac{\rho_{\rm m}}{\rho_{\rm crit}}=\frac{\rho_{\rm m,i}\left(a_{i}/a\right)^{3}}{\rho_{\rm crit,i}\left(H/H_{i}\right)^{2}}=\left(\frac{H_{i}}{H}\right)^{2}\Omega_{\rm m,i}\left(\frac{a_{i}}{a}\right)^{3}$
$\Omega_{\rm r}$ $=$ $\frac{\rho_{\rm r}}{\rho_{\rm crit}}=\frac{\rho_{\rm r,i}\left(a_{i}/a\right)^{4}}{\rho_{\rm crit,i}\left(H/H_{i}\right)^{2}}=\left(\frac{H_{i}}{H}\right)^{2}\Omega_{\rm r,i}\left(\frac{a_{i}}{a}\right)^{4}$
$\Omega_{\rm DE}$ $=$ $\frac{\rho_{\rm DE}}{\rho_{\rm crit}}=\frac{c_{H}\Lambda_{\rm QCD}^{3}H}{\frac{3H^{2}}{8\pi G}}=\frac{8\pi G}{3}\Lambda_{\rm QCD}^{3}\frac{c_{H}}{H}=\frac{\overline{H}}{H}$ (9)
Where $\overline{H}\equiv c_{H}\frac{8\pi G}{3}\Lambda_{\rm QCD}^{3}$ includes the coefficient $c_{H}$ to be consistent with (II.3) and (3). The $\overline{H}$ has the physical meaning of the Hubble constant at asymptotically far future when the Universe assumes an exact de Sitter state. The $(H_{i},a_{i})$ are specified at an arbitrary reference time $i$. We avoid using $a_{0}=1$ (i.e., anchoring the Friedmann equation at the present time) because of the dark energy term that is proportional to $H^{-1}$. Instead, we may prefer to anchor the solution at a time when the classical Friedmann equation holds, which means that $a_{0}$ is not necessarily equal to one today. The Friedmann equation can now be rewritten as:
$H^{2}-\overline{H}H-H_{\rm i}^{2}\left[\Omega_{\rm m,i}\left(\frac{a_{i}}{a}\right)^{3}+\Omega_{\rm r,i}\left(\frac{a_{i}}{a}\right)^{4}\right]=0$ (10)
This is a second order polynomial in $H$, which solution is given by:
$H(a)=\frac{\overline{H}}{2}\left(1+\sqrt{1+B\left(\frac{a_{i}}{a}\right)^{3}+C\left(\frac{a_{i}}{a}\right)^{4}}\right),$ (11)
where $B\equiv 4\left(\frac{H_{i}}{\overline{H}}\right)^{2}\Omega_{m,i}$ and $C\equiv 4\left(\frac{H_{i}}{\overline{H}}\right)^{2}\Omega_{r,i}$. It is straightforward to verify that Eq. (11) reduces to the classical Friedmann equation with only matter and radiation when $\overline{H}\rightarrow 0$ when $\sqrt{B}$ and $\sqrt{C}$ are both proportional to $\overline{H}^{-1}$ to cancel $\overline{H}$ in front. It is also easy to verify that $H(a\rightarrow\infty)\rightarrow\overline{H}$, which confirms the physical meaning of $\overline{H}$ as an asymptotic value for the Hubble constant in far future.
III.2 The de Sitter limit
We first want to compute the solution of Eq. (11) and the DE EoS when $a\gg 1$, i.e. when the system is not in the exact de Sitter state, but approaching it. In this limit, the radiation term can be neglected and we can chose the time $i$ to be today, such we can express the deviation from de Sitter as we approach today’s time from the future. In that case, we have $c_{H}$ is a constant and Eq. (11) can be written as:
$H=\frac{\overline{H}}{2}+\frac{\overline{H}}{2}\sqrt{1+4\Omega_{M,0}\left(\frac{H_{0}^{2}}{\overline{H}^{2}}\right)\left(\frac{a_{0}}{a(t)}\right)^{3}},$ (12)
In the de Sitter limit, when $a\rightarrow\infty$, the matter term can be neglected too, and we have $H\rightarrow\overline{H}$. Differentiating with respect to time $t$, we get:
$\dot{H}=\frac{\partial H}{\partial a}{H}a=-\frac{3\overline{H}H\Omega_{M,0}\left(\frac{a_{0}}{a(t)}\right)^{3}}{\sqrt{1+4\Omega_{M,0}\left(\frac{H_{0}^{2}}{\overline{H}^{2}}\right)\left(\frac{a_{0}}{a(t)}\right)^{3}}}$ (13)
where we use $\dot{a}={H}a$. Using the acceleration equation and the DE equation of state $P_{\rm DE}=w\rho_{\rm DE}$, we can write $\dot{H}$ as a sum over the mass densities:
$\dot{H}=-4\pi G\left(\rho_{m}+\rho_{\rm DE}(1+w)\right)$ (14)
Combining the DE mass density $\rho_{\rm DE}\equiv c_{H}\Lambda_{\rm QCD}^{3}H$ with Eq.(12), we get:
$\rho_{\rm DE}=\frac{3\overline{H}H}{8\pi G}=\frac{3\overline{H}^{2}}{16\pi G}\left[1+\sqrt{1+4x(t)}\right],$ (15)
where we have defined $x(t)$:
$x(t)$ $\equiv$ $\Omega_{M,0}\left(\frac{H_{0}^{2}}{\overline{H}^{2}}\right)\left(\frac{a_{0}}{a(t)}\right)^{3}$ (16)
$=$ $\Omega_{M,0}\left(\frac{H_{0}^{2}}{\overline{H}^{2}}\right)(1+z)^{3}.$~~
Similarly, we can express the matter density as:
$\rho_{\rm m}=\frac{3\overline{H}^{2}}{8\pi G}x(t)$ (17)
Combining Eqs. (14, 15, 17) we obtain the DE EoS parameter that is assymptotically justified for $x(t)\ll 1$, i.e. as long as DE is the dominant contributor to the Universe expansion:
$w+1=\frac{\rho_{\rm DE}+P_{\rm DE}}{\rho_{\rm DE}}=\frac{\left[\frac{2x(t)}{\sqrt{1+4x(t)}}\right]}{\left[1+\sqrt{1+4x(t)}\right]},$ (18)
One can clearly see that $w\rightarrow-1$ as $x(t)\rightarrow 0$, recovering our earlier Eq. (7) for the de Sitter state. Interestingly, as we approach the present time, when $x(t_{0})\sim 1$, Eq.(18) yields $w\simeq-0.7$, illustrating two important consequences of our framework, in agreement with the DESI results: (1) the equation-of-state parameter $w$ is time-dependent, and (2) it is indeed possible to have w>1w>-1 at the present epoch.
To what extend can Eq.(18) be used at the present day? A qualitative answer can be obtained from the requirement that the adiabatic approximation should be justified, i.e.
$|\dot{H}|\ll H^{2}.$ (19)
Substituting the corresponding expressions for $\dot{H}$ and $H^{2}$ in terms of $\rho_{\rm DE}$, $P_{\rm DE}$, and $\rho_{\rm M}$, the following condition must be satisfied:
$\left(1+\frac{w\rho_{\rm DE}}{\rho_{\rm DE}+\rho_{\rm M}}\right)\ll\frac{3}{2},$ (20)
which can be only justified when $\rho_{\rm M}\ll\rho_{\rm DE}$ in which case $w\approx-1$, and left hand side becomes indeed much smaller than right hand side of (20). The left hand side is $\approx 0.5$ today, such that (20) is marginally satisfied today at $z=0$.
III.3 Solution for $z\geq 0$
How can our formalism be implemented for $z\geq 0$, and specifically, how to solve Eq.(11) in the general case? As discussed in Section II, when the QCD induced DE mechanism operates with a constant $\overline{H}$, the system cannot deviate significantly from pure de Sitter geometry. If the condition (20) is not satisfied, the QCD-induced dark energy is not generated, since it was initially derived under the assumption of a pure de Sitter geometry, not a generic Friedmann Universe.
The simplest approach is to introduce a time dependent $\beta(t)$ which is approaching a constant value $\beta(z\rightarrow-1)\rightarrow 1$ in future when the de Sitter dominates the Universe evolution, while $\beta(z\gtrsim 1)\rightarrow 0$ vanishes at sufficiently large redshift. The DE density is now parameterized as follows:
$\rho_{DE}=\beta(t)c_{H}\Lambda_{\mathrm{QCD}}^{3}{H},~~~~~~~\beta(t)\in(0,1)$ (21)
The quantity $\beta(t)$ functions as a switch that activates dark energy at a certain redshift. It can be interpreted as a physically motivated parametrization of dark energy within our framework, where dark energy is induced by QCD. This contrasts with the canonical parametrization $w(a)=w_{0}+w_{a}(1-a)$, which merely captures a linear deviation from w=1w=-1. It is important to emphasize that introducing a time-dependent $\beta(t)$ is not an ad hoc addition to the proposal. Rather, it serves as an effective phenomenological tool to constrain the region of validity of our formulae near the de Sitter regime, where they were derived—hence its physical motivation. In principle, as previously mentioned, the numerical coefficient $\beta(t)c_{H}]$ can be calculated from first principles for any given geometry at a given time $t$, followed by the appropriate subtraction of the corresponding Minkowski space value, in analogy with the subtraction procedure performed in (II.3) for pure de Sitter space with constant $\overline{H}$.
Figure 1: Plot of the two functions $\beta(t)$ given in Eq.(24). $t_{i}=0.0465\;{\rm Gyr}$ corresponds to the anchor point with the Friedmann cosmology. it is very close to the Big-Bang so that $t-t_{i}=0$ can be considered to be the beginning of the expansion of the Universe. The grey area above $t-t_{i}=13.78\;{\rm Gyr}$ corresponds to the future.
[…]
Figure 2: Plots comparing various quantities in our framework to the Planck2018 cosmology as function of redshift. Red and black curves correspond respectively to the activation functions $\beta_{1}(t)$ and $\beta_{2}(t)$. When present, the green curve correspond to the Planck 2018 cosmology. Panel (a): fractional difference of the Hubble parameter, panel (b): fractional difference for the DE mass density $\rho_{\rm DE}$, panel (c): DE dimensionless density parameter, panel (d): DE equation of state parameter $w_{\rm DE}$. The grey area for $z<0$ corresponds to the future.
[…]
$\frac{da}{d\tau}=\beta(\tau)\frac{a}{2}\left(1+\sqrt{1+\frac{B}{\beta^{2}}\;\left(\frac{a_{i}}{a}\right)^{3}+\frac{C}{\beta^{2}}\;\left(\frac{a_{i}}{a}\right)^{4}}\right)$ (22)
and solve for $a(\tau)$. In expression (22) we explicitly separated the time dependent portion from coefficients $B,C$ such that $\overline{H}=c_{H}(\frac{8\pi G}{3}\Lambda_{\rm QCD}^{3})$ remains to be a constant as it assumes its de Sitter value at asymptotically far future when $\beta=1$, see Section III.1.
The boundary condition is set at $z_{i}=50$, using the cosmological parameters values from the Planck cosmology $\Omega_{\rm m,i}\simeq 0.982$ and $\Omega_{\rm r,i}\simeq 0.018$. The DE parameter at $z_{i}=50$ is $\Omega_{\rm DE,i}\simeq 1.65\times 10^{-5}\ll 1$, as required. At such high redshift, the Hubble parameter is $H_{i}\simeq 1.38\times 10^{5}\;{\rm km/s/Mpc}$, which implies that $B,C\gg 1$, without any fine-tuning of $\overline{H}$ and considering that $\beta(t)\ll 1$ at high redhsift. We use $a_{i}=1/(1+z_{i})\simeq 0.0196$ at time $\tau_{i}$, where $t_{i}\simeq 0.0465\;{\rm Gyr}$.
$\beta_{1}(t)$ $=$ $\frac{1}{2}\left[1+\operatorname{erf}\left(\frac{t-t_{1}}{\Delta t_{1}}\right)\right]\frac{1}{1+\exp\left(\frac{-(t-t_{2})}{\Delta t_{2}}\right)}$ (23)
$\beta_{2}(t)$ $=$ $\frac{1}{1+\exp\left(\frac{-(t-t_{3})}{\Delta t_{3}}\right)}$ (24)
where we set $t_{1}=\Delta t_{1}=5\;{\rm Gyr}$, $t_{2}=1\;{\rm Gyr}$, $\Delta t_{2}=0.5\;{\rm Gyr}$, $t_{3}=3\;{\rm Gyr}$, and $\Delta t_{3}=1\;{\rm Gyr}$. These values do not correspond to any specific physical event; they are selected solely to illustrate two different time scales at which the QCD-induced dark energy can become active in the Universe. The functions $\beta_{1}$ and $\beta_{2}$ are shown in Figure 1.
Equation (22) is then solved numerically with $\overline{H}=50\;{\rm km\;s^{-1}\;Mpc^{-1}}$. Figure 2 shows the variation of the Hubble parameter $H(z)$, the dark energy density $\rho_{\rm DE}(z)$, the density parameter $\Omega_{\rm DE}(z)$, and the dark energy equation of state $w_{\rm QCD}(z)$ as functions of redshift, and compares them to the predictions of the Planck cosmology. Panels (a), (b), and (c) illustrate that different histories of QCD dark energy activation can influence key cosmological quantities, which are fundamental for cosmological tests. Panel (d) shows that, within our framework, the dark energy equation of state parameter can vary with time, and may even cross the w=1w=-1 line, even multiple times. Although our choice of the $\beta(t)$ function is entirely arbitrary and not based on any physical model, it is still possible to test the self-consistency of the framework by performing all relevant cosmological tests. If a single $\beta(t)$ function proves successful across all these tests, it would provide strong evidence in support of our approach. The main result of this numerical experiment is that it is straightforward to obtain an equation of state parameter that varies with time and can take values both above and below 1-1 without the need for a new dynamical field.
Panel (d) also shows that for $z<0$ (i.e., in the future), $w_{\rm DE}$ converges to 1-1, consistent with the derivation in Section III.2. This is a generic feature of our framework: we predict that the de Sitter phase will be asymptotically reached in the future, implying that today we should observe $w_{\rm DE}>-1$. The commonly used parametrization $w(a)=w_{0}+w_{a}(1-a)$ does not represent a good model in our framework because it does not capture the phase towards de Sitter described in Section III.2. For $z>0$, however, $w_{\rm DE}$ can take any form as a function of redshift, depending on the activation function of the QCD-induced dark energy, $\beta(t)$. Although we cannot derive $\beta(t)$ theoretically, it is likely to be strongly constrained by observational data. Moreover, by comparing panels (a) and (d), one can see a strong correlation between the Hubble parameter $H(z)$ and $w(z)$. This correlation is unique to our framework because it is coming from the fact that $\rho_{\rm DE}(t)\propto\beta(t)H(t)$ in recent times when DE started to dominate the evolution of the universe. It offers an additional testable prediction of the framework. What is needed for this test are independent measurements of both $H(z)$ and $w(z)$. Future data releases from DESI and Stage IV surveys (such as Euclid and LSST) will improve the precision of $w(z)$. While current independent measurements of the Hubble parameter at high redshifts are still imprecise, there is hope for significant improvement in the near future.
III.4 Possible relation to the $H_{0}$ tension?
Figure 3: Commoving distance as a function redshift $z$ relative to the Planck cosmology. The red abd black curves correspond to the activation functions $\beta_{1}(t)$ and $\beta_{2}(t)$ in Eq.(24).
[…]
$\theta_{\star}\equiv\frac{r_{s}(z_{\star})}{D_{A}(z_{\star})},$ (25)
where $D_{A}(z_{\star})$ is the angular diameter distance at recombination and $r_{s}(z_{\star})$ is the sound horizon at $z_{\star}$. It is determined by the early-time integral of the sound speed $c_{s}(z)$:
$r_{s}(z_{\star})=\int_{z_{\star}}^{\infty}\frac{c_{s}(z)}{H(z)}dz$ (26)
On the other hand, BAO measurement take place at some late-time redshift $z_{\rm low}$. It measures the sound horizon angular scale $\theta_{d}(z_{\rm low})$:
$\theta_{d}\equiv\frac{r_{s}(z_{d})}{D_{A}(z_{\rm low})},$ (27)
with $r_{s}(z_{d})\simeq 1.018r_{s}(z_{\star})$ and where $z_{d}\sim 1059$ is the baryong drag redshift, which takes place slightly after recombination $z_{\star}\sim 1090$, when baryons are fully released from the photons drag. The impossibility of resolving the Hubble tension through a single early- or late-time modification arises from the fact that Eqs. (25) and (27) cannot be simultaneously satisfied under current cosmological constraints (Bernal et al., 2016).
It has been suggested that combining pre- and post-recombination new physics may offer a more satisfactory resolution of the $H_{0}$ tension, while simultaneously alleviating the amplitude-of-fluctuations tension in $S_{8}$ (Vagnozzi et al., 2022; Adil et al., 2023; Poulin et al., 2024). In our framework, the Friedmann equation is modified at all redshifts, with the detailed behavior determined by the unknown function $\beta(z)$. Using the two ad hoc choices for $\beta(z)$ introduced in Eq. (24), Figure 3 demonstrates that comoving distances can be altered across the entire redshift range. This behavior parallels the evolution of the Hubble constant shown in Panel (a) of Fig. 2. Therefore, an important open question is whether there exists a function $\beta(z)$ capable of satisfying all current cosmological constraints from CMB, BAO, SN, and LSS, how many free parameters such a function would require, and to what extent it could provide a physically viable description of the QCD vacuum. A full statistical analysis using a realistic parametrization of $\beta(z)$ is left to future work.
Note that, under the conjecture that Eq. (1) applies to the Hubble parameter, spatial variations in the topological sectors of the QCD vacuum may arise because local energy-density fluctuations would modify the Hubble rate. This opens the possibility of behavior resembling clustered dark energy, or more precisely, a dark-energy equation of state that correlates with the local mass density. Within our framework, such behavior can emerge without introducing any exotic additional fields, and may provide a new avenue for dark energy coupled to the dark sector (Pitrou and Uzan, 2024).
IV Conclusion
We have presented a physically motivated model of dark energy (DE) rooted in the topological structure of the QCD vacuum, as developed in earlier works Zhitnitsky (2014, 2015); Barvinsky and Zhitnitsky (2018). In this framework, DE is identified with the leftover vacuum energy generated by tunnelling transitions between different topological sectors, following the subtraction of a large contribution computed in Minkowski spacetime—an approach known as the Zeldovich prescription Zeldovich (1967). The amplitude of DE is determined by the QCD energy scale, as given in Eq. (3), while its time dependence arises from the evolution of the topological sectors $|k\rangle$ and the corresponding variation in tunnelling transition rates in an expanding universe.
This framework requires no new fields, coupling constants, or fine-tuning in the parameter estimates of Eq. (3) as everything is expressed in terms of a single QCD parameter, $\Lambda_{\mathrm{QCD}}$. It is founded entirely on the well-established physics of vacuum topology in non-Abelian gauge theories—an area whose cosmological implications have largely been overlooked. This stands in stark contrast to conventional dark energy models, which typically invoke a dynamical scalar field $\phi$ (e.g., quintessence, k-essence, phantom fields). Such models require an extremely flat potential, with a characteristic mass scale $m_{\phi}\simeq 10^{-33}$ eV—an unnaturally small value by particle physics standards, implying substantial fine-tuning. Furthermore, phantom field scenarios often suffer from serious theoretical issues, including quantum instabilities and violations of unitarity, which challenge their consistency with fundamental principles of quantum field theory.
The QCD-induced dark energy model offers a natural resolution to several long-standing fine-tuning problems, including the “coincidence problem,” the “drastic separation of scales,” and the “unnatural weakness of interactions.” In this framework, the dark energy scale is set by $\Lambda_{\mathrm{QCD}}$, placing it on the same order of magnitude as the observed value. Moreover, the question of “why now?” receives a compelling explanation: the timescale at which $\rho_{\rm DE}$ becomes dynamically relevant is estimated as $t_{0}=\overline{H}^{-1}\simeq 10$ Gyr. This timescale is also determined by $\Lambda_{\mathrm{QCD}}$, yielding the correct order of magnitude for the present-day Hubble scale.
[…]
1.In our framework, the Friedmann equation predicts that the dark energy equation of state parameter will asymptotically approach $w_{\rm DE}=-1$ in the future;
[…]
3. The QCD-induced dark energy leads to solutions of the Friedmann equation in which $w_{\rm DE}$ may lie above or below 1-1 and can cross the w=1w=-1 boundary multiple times throughout the history of the Universe, a trend suggested by the DESI results DESI Collaboration et al. (2025).
4. Notably, because $\rho_{\rm DE}\propto\beta H$ in near vicinity of the de Sitter stage when the DE dominates the evolution, our approach suggests a unique connection between $w_{\rm DE}(z)$ and the Hubble parameter $H(z)$, providing a distinctive, testable prediction. This relation may offer a natural explanation for the observed tension between the local measurement of the Hubble constant $H_{0}$ and its value inferred from observations at $z\sim 1100$, when the dark energy density was negligible.
5. The de Sitter regime described by the equation of state (7) would represent the final state of our Universe if the interaction between QCD gauge configurations (which saturate the vacuum energy) and massless electromagnetic photons were permanently switched off, as noted in item 1 above. However, when the coupling between the QCD vacuum fields and the electromagnetic field is restored, the departure from de Sitter behavior is triggered precisely by this interaction, which is unambiguously determined by the triangle anomaly, as discussed in Zhitnitsky (2019).
6. Finally, the QCD-induced dark energy framework predicts the generation of a cosmic magnetic field with an estimated strength on the order of $10^{-10}$ G, as argued in Zhitnitsky (2019). Remarkably, the corresponding correlation length spans the scale of the observable Universe—on the order of gigaparsecs. Intriguingly, observational evidence for magnetic fields with such enormous correlation lengths has indeed been reported, as reviewed in Durrer and Neronov (2013).
Interestingly, is it, at least in principle, possible to test some of the key ideas underlying this dark energy framework—specifically its origin in tunnelling processes between different topological sectors—using tabletop experiments? The ultimate answer is affirmative: this represents a genuine physical phenomenon rather than a mere formal reinterpretation of equations. The basic concept behind such an experiment is to detect a novel contribution to the Casimir vacuum energy in Maxwell theory, as proposed in Cao et al. (2013, 2016); Yao and Zhitnitsky (2017). This contribution, known as the Topological Casimir Effect, arises from non-dispersive tunnelling effects rather than from conventional quantum fluctuations of propagating photons with two physical transverse polarizations (dispersive contribution). Although this correction to the Casimir pressure has not yet been observed, its detection would provide direct empirical support for the topological origin of dark energy.
Our final remark concerning possible future developments of this work is as follows. It is well known that de Sitter–like behavior has occurred twice in the history of the Universe: first during the inflationary epoch, and again in the present epoch dominated by dark energy. The dark energy framework explored in this paper may offer valuable insights into the inflationary phase. Indeed, in a purely hypothetical scenario proposed in Zhitnitsky (2014); Barvinsky and Zhitnitsky (2018), the vacuum energy responsible for inflation could also arise from tunnelling transitions in a novel, as yet unidentified, strongly coupled gauge theory—an idea analogous to the QCD-based mechanism we advocate here for generating the dark energy scale, as described in Eq. (3). While the specific gauge theory that might play the role of QCD in this inflationary context remains unknown, the relevant energy scale (analogous to $\Lambda_{\mathrm{QCD}}$ for dark energy) and the expected number of e-foldings can nevertheless be estimated within this framework Zhitnitsky (2014); Barvinsky and Zhitnitsky (2018).
Acknowledgements
We are very thankful to Robert Brandenberger for the long and very useful discussions during his visit to UBC. In fact, we are grateful to Robert for the initiation and motivation of this project, and also for insisting to complete the writing in a timing manner. This research was supported in part by the Natural Sciences and Engineering Research Council of Canada.
Appendix A On use and misuse of the so-called Ghost DE model
The main goal of this Appendix is twofold. First, we aim to reiterate the key conceptual elements concerning the nature of dark energy as proposed in this work. Second, we seek to clarify the physical interpretation of the energy expression given in Eq. (6), in order to prevent potential misuse or misinterpretation of this formula.
The main features of dark energy in this framework were outlined in Section II.2. A key point is that, while dark energy may depend on time (through its dependence on $H$), this time variation cannot be described in terms of any local dynamical fields—a method typically employed and widely accepted in the community to introduce time dependence. As discussed above, the evolution of dark energy in this framework arises instead from modifications to the topological sectors $|k\rangle$ and the corresponding tunnelling transition rates in an expanding universe. This is precisely why the system remains sensitive to arbitrarily large distances, despite the presence of a mass gap, as explained in Section II.1.
The underlying physics can indeed be described using auxiliary topological fields, which offer an alternative to explicit non-perturbative calculations, as reviewed in Section II.2. These fields are non-dynamical, as they lack a canonically conjugate momentum. A concrete construction of such an auxiliary topological field has been carried out in the exactly solvable deformed QCD model Zhitnitsky (2013), where essential features such as topological sectors $|k\rangle$ and non-trivial holonomy are explicitly realized. In this context, it becomes evident that the auxiliary field introduced in Zhitnitsky (2013) does not propagate and possesses no kinetic term. Instead, it functions purely as a Lagrange multiplier within the system. Moreover, the energy described by Eq. (6) and generated by this auxiliary field cannot be attributed to any propagating degrees of freedom—it constitutes a so-called non-dispersive contribution.
It turns out that these topological fields exhibit precisely the properties of the Veneziano ghost Veneziano (1979), which was originally postulated to resolve the so-called $U(1)_{A}$ problem in QCD.  [^7] To our knowledge, the explicit construction of the Veneziano ghost in an exactly solvable model, as presented in Zhitnitsky (2013), remains the only known example where one can observe the development of a $1/q^{2}$ pole at zero momentum with a residue of the “wrong” sign—hence the designation as a ghost. The same construction also demonstrates that the Veneziano ghost is non-propagating, thereby avoiding any conflict with fundamental principles of quantum field theory such as unitarity and causality. This connection between dark energy and the Veneziano ghost, as originally proposed in Urban and Zhitnitsky (2010a, b), led to the model being named “Ghost Dark Energy” (GDE). A large body of literature has since explored various aspects and generalizations of this idea. However, many of these works have been accompanied by significant misconceptions regarding the theoretical foundations of GDE.
It is not our intention to provide a comprehensive critique of all the misconceptions and misinterpretations found in the literature. However, we do wish to highlight several common misunderstandings that can be particularly misleading when interpreting the expression for the vacuum energy given in Eq.(6). We emphasize these points in order to prevent misinterpretation of our results on dark energy, which are briefly summarized in SectionIII and are also based on Eq. (6).
In particular, in Cai et al. (2011) and many subsequent papers, the expression given in Eq. (6) was interpreted as if it were generated by a dynamical field $\phi$, with the equation of state (EoS) and the speed of sound $c_{s}$ expressed in terms of this field $\phi$ as follows:
$w=\frac{\dot{\phi}^{2}-2V(\phi)}{\dot{\phi}^{2}+2V(\phi)},~~~~~~~c_{s}^{2}\equiv\frac{{\dot{p}_{DE}}}{{\dot{\rho}_{DE}}}.$ (A1)
From the expressions for $w$ and $c_{s}$, it becomes evident that the system exhibits serious issues—such as violations of unitarity, causality, and stability—when w<1w<-1 or $c_{s}^{2}<0$. However, these problems arise only when the underlying model assumes a dynamical field. In contrast, when the variation originates from an auxiliary (non-dynamical) field, as in our framework, such behavior does not conflict with any fundamental principles, since there are no propagating degrees of freedom—there is simply nothing to propagate. Consequently, concerns about classical stability for cases where $c_{s}^{2}<0$, as discussed for example in Ebrahimi and Sheykhi (2011); Biswas et al. (2019), are not relevant to this model. The computation of a speed of sound in this context is meaningless, as there is no physical field to support fluctuations.
Another common misconception is as follows. If one interprets the behavior with w<1w<-1—which is precisely what we have found in Section III—as indicative of phantom dark energy, then, following Eq. (A1), one could reconstruct a specific scalar potential $V(\phi)$ that would yield such an equation of state, as discussed in Das and Bharali (2021). However, as we have emphasized, there are no propagating degrees of freedom in our framework. The occurrence of w<1w<-1 is entirely consistent with the fundamental principles of quantum field theory, precisely because it does not involve any dynamical field. Consequently, within our model, describing a regime with w<1w<-1 does not necessitate the introduction of a fundamental phantom field, in stark contrast to the conventional interpretation.
To conclude this Appendix, which has addressed common misinterpretations of Eq. (6) and its various implications—including the regime with w<1w<-1—we emphasize that the so-called phantom behavior in our framework is fully consistent with the fundamental principles of quantum field theory. This consistency arises from the fact that the underlying dynamics is not associated with any propagating or dynamical degrees of freedom.
Appendix B Few technical comments on the nature of the QCD-induced DE
Below, we briefly review several technical elements that were discussed in the main text in Section II.2:
1. The main claim of Zhitnitsky (2015) is the presence of a linear correction, as expressed in Eq.(1). This constitutes a central element of the entire proposal, as one would naively expect a quadratic correction at small curvature, i.e., $\propto\kappa^{2}$. Indeed, the only difference between the two geometries $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{R}^{3}\times\mathbb{S}^{1}$—assuming identical sizes for $\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{S}^{1}$—is the small curvature $\sim\kappa^{2}$ of the hyperbolic space $\mathbb{H}^{3}_{\kappa}$, in contrast to the zero curvature of $\mathbb{R}^{3}$. Conventional locality arguments would therefore unambiguously suggest that any correction to the vacuum energy difference in Eq.(1) should be expressible solely in terms of the curvature, and thus involve only even powers of $\kappa$, i.e., $\propto\kappa^{2n}$. However, explicit calculations in Zhitnitsky (2015) reveal a linear correction, contradicting this expectation.
2. This linear correction, $\sim\kappa$, which plays a central role in the present proposal, is generated by vacuum configurations with nontrivial holonomy. The holonomy is defined as $U(\mathbf{x})\equiv{\cal{P}}\exp\left(i\int_{0}^{{\cal{T}}}dx_{4}A_{4}(x_{4},\mathbf{x})\right)$, where $\mathcal{P}$ denotes path ordering. Specifically, we consider the Polyakov line evaluated at spatial infinity, i.e.
$L={\cal{P}}\exp\left(i\int_{0}^{{\cal{T}}}dx_{4}A_{4}(x_{4},|\mathbf{x}|\rightarrow\infty)\right).$ (B1)
The operator $\mathrm{Tr}L$ classifies the self-dual solutions that may contribute to the path integral at finite temperature, $T\equiv\mathcal{T}^{-1}$, including the low-temperature limit as $T\rightarrow 0$. In particular, for $SU(2)$ gauge group the holonomy
$\frac{1}{2}\mathrm{Tr}L=\cos(\pi\nu),$ (B2)
belongs to the group center $\frac{1}{2}\mathrm{Tr}L=\pm 1$ when $\nu$ assumes the integer values (trivial holonomy). The so-called confining value for the holonomy corresponds to $\nu=1/2$ when $\mathrm{Tr}L=0$ vanishes.
Therefore, the standard arguments based on locality are strongly violated by such configurations. As shown in the computations of Zhitnitsky (2015), the linear correction $\sim\kappa$ is explicitly proportional to the holonomy defined in Eq. (B2), a gauge-invariant observable that cannot be reduced to the local curvature. In other words, this correction arises from non-local configurations and cannot be expressed in terms of the local curvature $\sim\kappa^{2}$.
3. The linear correction $\sim\kappa$ in $\Delta E_{\rm vac}$ can be traced, at a technical level, to the differing behavior of monopole configurations at large distances in the two backgrounds. In hyperbolic space $\mathbb{H}^{3}_{\kappa}$, monopole fields experience an exponential suppression at the characteristic scale $\kappa^{-1}$, whereas no such cutoff is present in flat space $\mathbb{R}^{3}$ Zhitnitsky (2015).
4. Another technical reason that makes these computations feasible is the conformal equivalence between $\mathbb{H}^{3}_{\kappa}\times\mathbb{S}^{1}_{\kappa^{-1}}$ and $\mathbb{R}^{4}$, which ensures that the subtraction procedure in Eq. (1) is well defined Zhitnitsky (2015). This is a crucial point because the number of zero modes (i.e., the moduli space) contributing to the path integral is identical in both geometries, and the expressions for their volumes coincide in the small $\kappa$ limit, given a fixed large radius. As a result, the computation yields an energy expression that is infrared finite and exhibits extensive scaling with volume, i.e., $E\propto\rho V$. This outcome is highly nontrivial in the context of non-perturbative calculations, where achieving extensivity typically requires summing over an infinite number of monopole configurations. For further details, see Zhitnitsky (2015).
5. One additional technical comment is as follows. As explained in Section II.2, the driving mechanism behind dark energy in this framework is the tunnelling between QCD topological sectors, rather than the dynamics of a real propagating scalar field $\phi$. This fundamental distinction in the nature of dark energy can be reformulated using an auxiliary field—the so-called Veneziano ghost. For further references and details, see Appendix A. An important point to note is that tunnelling events always contribute to correlation functions with a sign opposite to that of a real dynamical scalar field $\phi$ (such as those appearing in quintessence or k-essence models). This is precisely why the auxiliary field that effectively encodes these tunnelling processes is referred to as a “ghost” field. Crucially, this field does not propagate and does not violate any fundamental principles of quantum field theory, as clarified in Appendix A. Instead, this auxiliary field generates the so called non-dispersive contact term, which cannot be expressed in terms of any physical propagating degrees of freedom (which, by definition, may generate only dispersive contributions).
6. This non-dispersive contact term cannot be removed by any UV renormalization procedures as it has the IR nature, see Barvinsky and Zhitnitsky (2018) for more details.